# Inverse Reinforcement Learning by Estimating Expertise of Demonstrators

### v1

@@ Abstract
In Imitation Learning (IL), utilizing suboptimal and heterogeneous demonstrations presents a substantial challenge due to the varied nature of real-world data. However, standard IL algorithms consider these datasets as homogeneous, thereby inheriting the deficiencies of suboptimal demonstrators. Previous approaches to this issue typically rely on impractical assumptions like high-quality data subsets, confidence rankings, or explicit environmental knowledge. This paper introduces IRLEED, Inverse Reinforcement Learning by Estimating Expertise of Demonstrators, a novel framework that overcomes these hurdles without prior knowledge of demonstrator expertise. IRLEED enhances existing Inverse Reinforcement Learning (IRL) algorithms by combining a general model for demonstrator suboptimality to address reward bias and action variance, with a Maximum Entropy IRL framework to efficiently derive the optimal policy from diverse, suboptimal demonstrations. Experiments in both online and offline IL settings, with simulated and human-generated data, demonstrate IRLEED’s adaptability and effectiveness, making it a versatile solution for learning from suboptimal demonstrations.
Learning from Suboptimal Demonstrations, Inverse Reinforcement Learning
Figure 1: IRLEED can be applied in the suboptimal demonstration setting to estimate the ground truth reward $r$, which is used to find the optimal policy. Left: A heterogeneous dataset is collected from multiple sources with varying degrees of optimality. We categorize this optimality by using accuracy to represent the reward bias, and precision to represent the variance in action choices. Right: We can infer the demonstrator policies using a model for suboptimal behaviour based on the Boltzmann rationality principle, which captures both the accuracy $\epsilon_{i}$, and the precision $\beta_{i}$, as compared to the ground truth reward $r$. Middle: Using estimates of the demonstrator policies $\hat{\pi}_{i}$ along with the demonstrations $\mathcal{D}_{i}$, we can optimize for the true reward $r$, as well as the parameters that capture accuracy $\epsilon$ and precision $\beta$.
@@ 1 Introduction
Reinforcement Learning (RL) has proven to be a versatile and powerful tool across a wide range of applications, from navigating autonomous vehicles, to playing complex games like Go and Chess. The efficacy of RL algorithms is largely attributed to their ability to optimize hand-crafted reward functions (Brockman et al., 2016). However, designing these functions is often challenging and impractical in complex environments (Hadfield-Menell et al., 2017). A prevalent alternative to this dilemma is Imitation Learning (IL), which bypasses the need for cumbersome reward engineering by leveraging expert demonstrations to instill desired behaviors (Argall et al., 2009).
@@
The two main approaches utilized in IL are: behavioral cloning (Pomerleau, 1991), which acquires a policy through a supervised learning approach using state-action pairs provided by the expert, and Inverse Reinforcement Learning (IRL) (Ng & Russell, 2000), which solves for the reward function that makes the expert’s behavior optimal, subsequently facilitating the training of an IL policy. Despite the effectiveness of behavior cloning in simple environments with large amounts of data, learning a policy to fit single time step decisions leads to compounding errors due to covariance shift (Ross & Bagnell, 2010). On the other hand, IRL learns a reward function that prioritizes entire trajectories over others, considering the sequential nature of the decision making problem at hand (Abbeel & Ng, 2004). Consequently, the success of IRL has spurred the development of IL techniques that either explicitly or implicitly incorporate environment dynamics (Ho & Ermon, 2016; Fu et al., 2018; Kostrikov et al., 2019; Garg et al., 2021).
@@
However, a critical assumption in these methodologies is the availability of high-quality demonstration data. In many practical situations, especially with crowd-sourced or varied data, the quality of demonstrations is inconsistent (Mandlekar et al., 2021; Belkhale et al., 2023). For example, foundation models like GPT-3 which are trained on large and uncurated datasets, exhibit limitations such as generating biased outputs that negatively stereotype certain groups (Abid et al., 2021). In robotics, data curation becomes vital, since utilizing smaller real world datasets with noise or biases can lead to hazardous situations. Furthermore, assuming uniformity in demonstration quality overlooks the unique intentions of individual demonstrators, potentially leading to suboptimal learning outcomes (Eysenbach et al., 2018). Therefore, it is critical to develop IL methods that can account for the heterogeneity and suboptimality of demonstration data.
@@
Addressing the challenges posed by learning from suboptimal and heterogeneous demonstrations is complex. One line of research focuses on heterogeneous demonstrations without considering their quality (Chen et al., 2020, 2023), assuming that all provided demonstrations are optimal. Other works consider suboptimal demonstrations, but require unrealistic prerequisites such as explicit knowledge of environment dynamics (Brown et al., 2020; Chen et al., 2021), a set of confidence scores over the demonstrations (Zhang et al., 2021; Cao & Sadigh, 2021; Wu et al., 2019; Kuhar et al., 2023), or direct access to a subset of expert demonstrations (Xu et al., 2022; Yang et al., 2021). A recently proposed approach that avoids these assumptions is ILEED (Beliaev et al., 2022), which directly models the expertise of demonstrators within a supervised learning framework, thus enabling one to learn from the optimal behavior while filtering out the suboptimal behavior of each demonstrator. While this method was proven effective, it is based on a behavioral cloning formulation, which overlooks the underlying environment dynamics, and its representation of suboptimal behavior is limited, presuming demonstrators to be noisy approximations of experts.
@@
To overcome these challenges, this paper introduces IRLEED, Inverse Reinforcement Learning by Estimating Expertise of Demonstrators. As illustrated in Figure 1, IRLEED is a novel framework for IRL that accounts for demonstrator suboptimality without prior expertise knowledge. It comprises two core components: (1) a general model of demonstrator suboptimality based on the Boltzmann rationality principle, which captures both the reward bias and the variance in action choices of each demonstrator as compared to the optimal policy (2) a Maximum Entropy IRL framework, which is used to efficiently find the optimal policy when provided with a set of suboptimal and heterogeneous demonstrations. These elements enable IRLEED to effectively recover the ground truth policy from suboptimal and heterogeneous demonstrations, surpassing the limitations of previous models. Furthermore, IRLEED’s simplicity and compatibility with existing IRL techniques make it a versatile and powerful tool in the field of IL.
@@
  • We empirically validate the success of our method against relevant baselines in both online and offline IL scenarios, utilizing both simulated and human datasets.
@@ 2 Preliminary
Notations We consider the standard Markov decision process (MDP) setting, defined by a tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{T},p_{0},r,\gamma)$, where $\mathcal{S}$, $\mathcal{A}$ represent state and action spaces, $\mathcal{T}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ represents the dynamics, $p_{0}:\mathcal{S}\rightarrow[0,1]$ represents the initial state distribution, $r\in\mathcal{R}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ represents the reward function, and $\gamma\in(0,1)$ represents the discount factor. In Section 3 and Section 4, we describe our method using finite state and action spaces, $\mathcal{S}$ and $\mathcal{A}$, but our experiments later utilize continuous environments. We use $\pi:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ to denote a policy that assigns probabilities to actions in $\mathcal{A}$ given states in $\mathcal{S}$. Considering the $\gamma$–discounted infinite time horizon setting, we define the expected value under a policy $\pi$ in terms of the trajectory it produces. Specifically, this expectation is represented as $\bar{f}_{\pi}=\operatorname{\mathbb{E}}_{\pi}[f(s,a)]$, which equals $\operatorname{\mathbb{E}}[\sum_{t=0}^{\infty}\gamma^{t}f(s_{t},a_{t})]$, where $s_{0}\sim p_{0},a_{t}\sim\pi(\cdot|s_{t})$, and $s_{t+1}\sim\mathcal{T}(\cdot|s_{t},a_{t})$. Additionally, we will use $\tilde{f}_{\mathcal{D}}=\operatorname{\mathbb{E}}_{\mathcal{D}}[f(s,a)]$ to denote the empirical expectation with respect to the trajectories $\tau_{j}=(s_{0},a_{0},\ldots,s_{T},a_{T})$ in demonstration set $\mathcal{D}=\{\tau_{j}\}_{j=1}^{M}$, where we leave out denoting trajectories of varying lengths for simplicity. Maximum Entropy IRL The goal of Inverse Reinforcement Learning (IRL) is to recover a reward $r\in\mathcal{R}$ that rationalizes the demonstrated behavior in dataset $\mathcal{D}$, where the provided dataset consists of trajectories $\tau$ sampled from an expert policy $\pi_{E}$. This general idea was first formulated as a feature-expectation matching problem (Abbeel & Ng, 2004): assuming that the feature vector $f:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}^{k}$ components quantify the expert’s behavior, to find a policy $\pi$ that performs equal to or better than the expert $\pi_{E}$, it suffices that their feature expectations match $\bar{f}_{\pi}=\bar{f}_{\pi_{E}}$. Unfortunately, this leads to an ill-posed problem as many policies can lead to the same feature counts. Maximum Entropy IRL resolves this ambiguity by choosing the policy which does not exhibit any additional preferences beyond matching feature expectations (Ziebart et al., 2008):
@@
where in the infinite-time horizon setting, this equates to maximizing the discounted causal entropy $H(\pi)$ under the feature matching constraint (Bloem & Bambos, 2014).
@@
with the parameterized policy $\pi_{\theta}$ defined recursively as:
[…]
The parameters $\theta\in\mathbb{R}^{k}$ specifying the soft Bellman policy, $\pi_{\theta}(a_{t}|s_{t})$, correspond to the dual variables of the feature matching constraint in Eq. (1). Without loss of generality, if we define the true reward signal as $r(s,a)=\theta^{\star\top}f(s,a)$ for some reward parameter ${\theta}^{\star}\in\mathbb{R}^{k}$ and features $f$, then the soft Bellman policy parameterized by $\theta^{\star}$ achieves the maximum possible return (for more details see (Ziebart, 2010), Theorem 6.2 and Corollary 6.11). [^1] To allow an arbitrary reward specification, the IRL objective defined in Eq. 1 can be expressed in terms of occupancy measures instead of feature expectations (Ho & Ermon, 2016). This formulates an iterative approach to finding the true reward parameter $\theta^{\star}$: given demonstration set $\mathcal{D}$ and your estimate $\theta$, alternate between inferring the policy $\pi_{\theta}(a_{t}|s_{t})$ based on Eq. (3), and optimizing the parameter $\theta$ based on Eq. (2). This line of work provides many insights that have been used to derive modern IRL (Fu et al., 2018; Reddy et al., 2019; Gao et al., 2018; Garg et al., 2021) and RL (Haarnoja et al., 2017, 2018) algorithms. Unlike these prior works, we consider the setting where we no longer have access to the expert policy $\pi_{E}$, and instead are given a heterogeneous set of potentially suboptimal demonstrations from which we want to learn a well performing policy. To accomplish this, we first describe the model we use to capture the suboptimality of demonstrators in Section 3, following which we frame this problem within the IRL framework and derive our own method IRLEED in Section 4. Finally, we describe our experimental results in Section 5, and conclude our work in Section 6.
@@ 3 Suboptimal Demonstrator Model
In this work, we consider the case where dataset $\mathcal{D}=\{(i,\mathcal{D}_{i})\}_{i=1}^{N}$ consists of a mixture of $N$ demonstrations with varying quality. Each demonstration $\mathcal{D}_{i}=\{\tau_{j}\}_{j=1}^{M_{i}}$ contains a set of $M_{i}$ trajectories sampled from a distinct, fixed policy $\pi_{i}$. Rather than considering $\mathcal{D}$ as a homogeneous set derived from one expert policy $\pi_{E}$, our approach focuses on leveraging the specific source of each demonstration to enhance the IRL framework. By identifying the varying qualities within this mixed collection of demonstrations, we aim to more accurately deduce the ground truth policy $\pi_{E}$, as opposed to a simplistic method that averages behaviors. To achieve this, it is essential to develop a model for the demonstrators $\pi_{i}$ that accurately reflects their suboptimal behavior relative to the optimal policy $\pi_{E}$.
A popular framework for modeling decision making is the Boltzmann rationality model, which is widely used in psychology (Baker et al., 2009), economics (Luce, 1959) and RL (Laidlaw & Dragan, 2021). In this model, the likelihood of an agent choosing a particular action is proportional to the exponential of the reward associated with that action: $\pi(a|s;r,\beta)\propto\exp(\beta r(s,a))$, where the inverse temperature parameter $\beta$ controls the randomness of the choice. While this formulation provides an intuitive way to capture one form of suboptimality through $\beta$, it traditionally assumes that agents have unbiased knowledge about the true reward $r(s,a)$. To broaden this model for suboptimal demonstrations, we propose considering agents that follow a biased reward $r(s,a)+\epsilon(s,a)$, introducing $\epsilon$ as the deviation from the true reward. This modification enables us to represent an agent’s accuracy through the deviation $\epsilon$, adding a bias to their perceived rewards, and their precision through parameter $\beta$, adding variance to their action choices.
Within the context of the IRL framework described in Section 2, we recall that the expert policy $\pi_{E}$ aligns with the true reward, formulated as $r(s,a)=\theta^{\star\top}f(s,a)$. Applying the aforementioned concept, we can model a demonstrator $i$’s perceived reward as a deviation from the true reward: $r_{i}(s,a)=(\theta^{\star}+\epsilon_{i})^{\top}f(s,a)$, where $\epsilon_{i}\in\mathbb{R}^{k}$. By incorporating this altered reward $r_{i}$ into the soft Bellman policy defined in Eq. (3), we derive a recursive parameterization for demonstrator $i$’s policy:
[…]
@@
Before moving forward, it is important to note that this model extends outside the scope of problems where the true reward $r$ is an affine transformation of some feature vector $f$. In the general setting, we can parameterize demonstrator $i$’s reward as $r_{i}=r+\epsilon_{i}$, a combination of the true reward $r\in\mathcal{R}$ and their deviation $\epsilon_{i}\in\mathcal{R}$, where $r$ and $\epsilon_{i}$ can be represented by neural networks to allow for general function approximation. We opt to formulate our proposed model within the classic feature matching scope, as this facilitates direct comparison with the seminal works used to derive our results in the following section. However, our proposed model’s applicability extends beyond this traditional framework, as we elaborate in Section 4.3, and demonstrate through experiments in Section 5. Until then, we utilize the feature matching framework described, referencing it as standard IRL.

### v2

@@ Abstract
In Imitation Learning (IL), utilizing suboptimal and heterogeneous demonstrations presents a substantial challenge due to the varied nature of real-world data. However, standard IL algorithms consider these datasets as homogeneous, thereby inheriting the deficiencies of suboptimal demonstrators. Previous approaches to this issue rely on impractical assumptions like high-quality data subsets, confidence rankings, or explicit environmental knowledge. This paper introduces IRLEED, Inverse Reinforcement Learning by Estimating Expertise of Demonstrators, a novel framework that overcomes these hurdles without prior knowledge of demonstrator expertise. IRLEED enhances existing Inverse Reinforcement Learning (IRL) algorithms by combining a general model for demonstrator suboptimality to address reward bias and action variance, with a Maximum Entropy IRL framework to efficiently derive the optimal policy from diverse, suboptimal demonstrations. Experiments in both online and offline IL settings, with simulated and human-generated data, demonstrate IRLEED’s adaptability and effectiveness, making it a versatile solution for learning from suboptimal demonstrations.
Codehttps://github.com/mbeliaev1/IRLEED
Extended versionhttps://arxiv.org/abs/2402.01886
Figure 1: IRLEED is applied in the suboptimal setting to estimate the ground truth reward $r$, which is used to find the optimal policy. Left: A heterogeneous dataset is collected from multiple sources with varying optimality. We categorize this optimality by using accuracy to represent the reward bias, and precision to represent the variance in action choices. Right: We infer the demonstrator policies using a model for behavior based on the Boltzmann rationality principle, which captures both the accuracy $\epsilon_{i}$, and the precision $\beta_{i}$, as compared to the ground truth reward $r$. Middle: Using estimates of the demonstrator policies $\hat{\pi}_{i}$ along with the demonstrations $\mathcal{D}_{i}$, we can optimize for the true reward $r$, and the parameters that capture accuracy $\epsilon$ and precision $\beta$.
@@ 1 Introduction
Reinforcement Learning (RL) has proven to be a powerful tool across a wide range of applications, from controlling robotic systems, to playing complex games like Go and Chess. The efficacy of RL algorithms is largely attributed to their ability to optimize hand-crafted reward functions (Brockman et al. 2016). However, designing these functions is often challenging and impractical in complex environments (Hadfield-Menell et al. 2017). A prevalent alternative to this dilemma is Imitation Learning (IL), which bypasses the need for cumbersome reward engineering by leveraging expert demonstrations to instill desired behaviors (Argall et al. 2009).
@@
The two main approaches utilized in IL are: behavioral cloning (Pomerleau 1991), which acquires a policy through a supervised learning approach using state-action pairs provided by the expert, and Inverse Reinforcement Learning (IRL) (Ng and Russell 2000), which solves for the reward function that makes the expert’s behavior optimal, subsequently facilitating the training of an IL policy. Despite the effectiveness of behavior cloning in simple environments with large amounts of data, learning a policy to fit single time step decisions leads to compounding errors due to covariance shift (Ross and Bagnell 2010). On the other hand, IRL learns a reward function that prioritizes entire trajectories over others, considering the sequential nature of the decision making problem at hand (Abbeel and Ng 2004). Consequently, the success of IRL has spurred the development of IL techniques that either explicitly or implicitly incorporate environment dynamics (Ho and Ermon 2016; Fu, Luo, and Levine 2018; Kostrikov, Nachum, and Tompson 2019; Garg et al. 2021).
@@
However, a critical assumption in these methodologies is the availability of high-quality demonstration data. In many practical situations, especially with crowd-sourced or varied data, the quality of demonstrations is inconsistent (Mandlekar et al. 2021; Belkhale, Cui, and Sadigh 2023). For robotics, data curation becomes vital, since utilizing smaller real world datasets with noise or biases can lead to hazardous situations. Furthermore, assuming uniformity in demonstration quality overlooks the unique intentions of individual demonstrators, potentially leading to suboptimal learning outcomes (Eysenbach et al. 2018). Therefore, it is critical to develop IL methods that can account for the heterogeneity and suboptimality of demonstration data.
@@
Related Work Addressing the challenges posed by learning from suboptimal and heterogeneous demonstrations is complex. One line of research focuses on heterogeneous demonstrations without considering their quality (Chen et al. 2020, 2023), assuming that all demonstrations are optimal. Other works consider suboptimal demonstrations, but require unrealistic prerequisites such as explicit knowledge of environment dynamics (Brown, Goo, and Niekum 2020; Chen, Paleja, and Gombolay 2021), a set of confidence scores over the demonstrations (Zhang et al. 2021; Cao and Sadigh 2021; Wu et al. 2019; Kuhar et al. 2023), or direct access to a subset of expert demonstrations (Shiarlis, Messias, and Whiteson 2016; Xu et al. 2022; Yang, Levine, and Nachum 2021).
A recent approach that avoids these assumptions is ILEED (Beliaev et al. 2022), which directly models the expertise of demonstrators within a supervised learning framework, thus enabling one to learn from the optimal behavior while filtering out the suboptimal behavior of each demonstrator. While this method was proven effective, it is based on a behavioral cloning formulation, which overlooks the underlying environment dynamics, and its representation of suboptimal behavior is limited, presuming demonstrators to be noisy approximations of experts.
@@
Overview To overcome these challenges, this paper introduces IRLEED, Inverse Reinforcement Learning by Estimating Expertise of Demonstrators. As illustrated in Figure 1, IRLEED is a novel framework for IRL that accounts for demonstrator suboptimality without prior expertise knowledge. It comprises two core components: (1) a general model of demonstrator suboptimality based on the Boltzmann rationality principle, which captures both the reward bias and the variance in action choices of each demonstrator as compared to the optimal policy (2) a Maximum Entropy IRL framework, which is used to efficiently find the optimal policy when provided with a set of suboptimal and heterogeneous demonstrations. These elements enable IRLEED to effectively recover the ground truth policy from suboptimal and heterogeneous demonstrations, surpassing the limitations of previous models. Furthermore, IRLEED’s simplicity and compatibility with existing IRL techniques make it a versatile and powerful tool in the field of IL.
@@
  • We empirically validate the success of our method against relevant baselines in both online and offline IL scenarios.
@@ 2 Preliminary
Notations We use the Markov decision process (MDP) setting, defined by a tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{T},p_{0},r,\gamma)$, where $\mathcal{S}$, $\mathcal{A}$ represent state and action spaces, $\mathcal{T}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ represents the dynamics, $p_{0}:\mathcal{S}\rightarrow[0,1]$ represents the initial state distribution, $r\in\mathcal{R}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ represents the reward function, and $\gamma\in(0,1)$ represents the discount factor. In Sections  3 and 4, we describe our method using finite state and action spaces, $\mathcal{S}$ and $\mathcal{A}$, but our experiments later utilize continuous environments. We use $\pi:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ to denote a policy that assigns probabilities to actions in $\mathcal{A}$ given states in $\mathcal{S}$. Considering the $\gamma$–discounted infinite time horizon setting, we define the expected value under a policy $\pi$ in terms of the trajectory it produces. Specifically, this expectation is represented as $\bar{f}_{\pi}=\operatorname{\mathbb{E}}_{\pi}[f(s,a)]$, which equals $\operatorname{\mathbb{E}}[\sum_{t=0}^{\infty}\gamma^{t}f(s_{t},a_{t})]$, where $s_{0}\sim p_{0},a_{t}\sim\pi(\cdot|s_{t})$, and $s_{t+1}\sim\mathcal{T}(\cdot|s_{t},a_{t})$. Additionally, we will use $\tilde{f}_{\mathcal{D}}=\operatorname{\mathbb{E}}_{\mathcal{D}}[f(s,a)]$ to denote the empirical expectation with respect to the trajectories $\tau_{j}=(s_{0},a_{0},\ldots,s_{T},a_{T})$ in demonstration set $\mathcal{D}=\{\tau_{j}\}_{j=1}^{M}$, where we leave out denoting trajectories of varying lengths for simplicity.
Maximum Entropy IRL The goal of IRL is to recover a reward $r\in\mathcal{R}$ that rationalizes the demonstrated behavior in dataset $\mathcal{D}$, where the provided dataset consists of trajectories $\tau$ sampled from an expert policy $\pi_{E}$. This idea was first formulated as a feature-expectation matching problem (Abbeel and Ng 2004): assuming that the feature vector $f:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}^{k}$ components quantify the expert’s behavior, to find a policy $\pi$ that performs equal to or better than the expert $\pi_{E}$, it suffices that their feature expectations match $\bar{f}_{\pi}=\bar{f}_{\pi_{E}}$. Unfortunately, this leads to an ill-posed problem as many policies can lead to the same feature counts. Maximum Entropy IRL resolves this ambiguity by choosing the policy which does not exhibit any additional preferences beyond matching feature expectations (Ziebart et al. 2008):
@@
where in the infinite-time horizon setting, this equates to maximizing the discounted causal entropy $H(\pi)$ under the feature matching constraint (Bloem and Bambos 2014).
@@
with the parameterized policy $\pi_{\theta}$ defined recursively using $Q^{\textit{soft}}$ and $V^{\textit{soft}}$:
$\pi_{\theta}(a_{t}|s_{t})$ $=\exp(Q^{\textit{soft}}_{\theta}(s_{t},a_{t})-V^{\textit{soft}}_{\theta}(s_{t})),$ (3)
$Q^{\textit{soft}}_{\theta}$ $=\theta^{\top}f(s_{t},a_{t})+\gamma\operatorname{\mathbb{E}}_{s_{t+1}\sim\mathcal{T}(\cdot|s_{t},a_{t})}[V^{\textit{soft}}_{\theta}(s_{t+1})],$ (4)
$V^{\textit{soft}}_{\theta}$ $=\log\sum_{a_{t}\in\mathcal{A}}\exp(Q^{\textit{soft}}_{\theta}(s_{t},a_{t})).$ (5)
The parameters $\theta\in\mathbb{R}^{k}$ specifying the soft Bellman policy, $\pi_{\theta}(a_{t}|s_{t})$, correspond to the dual variables of the feature matching constraint in Eq. (1). Without loss of generality, if we define the true reward signal as $r(s,a)=\theta^{\star\top}f(s,a)$ for some reward parameter ${\theta}^{\star}\in\mathbb{R}^{k}$ and features $f$, then the soft Bellman policy parameterized by $\theta^{\star}$ achieves the maximum possible return (for more details see (Ziebart 2010), Theorem 6.2 and Corollary 6.11). [^1] To allow an arbitrary reward specification, the IRL objective defined in Eq. 1 can be expressed in terms of occupancy measures instead of feature expectations (Ho and Ermon 2016). This formulates an iterative approach to finding the true reward parameter $\theta^{\star}$: given demonstration set $\mathcal{D}$ and your estimate $\theta$, alternate between inferring the policy $\pi_{\theta}(a_{t}|s_{t})$ based on Eq. (3), and optimizing the parameter $\theta$ based on Eq. (2).
This line of work provides many insights that have been used to derive modern IRL (Fu, Luo, and Levine 2018; Reddy, Dragan, and Levine 2019; Gao et al. 2018; Garg et al. 2021) and RL (Haarnoja et al. 2017, 2018) algorithms. Unlike these prior works, we consider the setting where we no longer have access to the expert policy $\pi_{E}$, and instead are given a heterogeneous set of potentially suboptimal demonstrations from which we want to learn a well performing policy. To this end, we first describe the model we use to capture the suboptimality of demonstrators in Section 3, following which we frame this problem within the IRL framework and derive our own method IRLEED in Section 4. Finally, we describe our experimental results in Section 5, and conclude our work in Section 6.
@@ 3 Suboptimal Demonstrator Model
This work considers the case where dataset $\mathcal{D}=\{(i,\mathcal{D}_{i})\}_{i=1}^{N}$ consists of a mixture of $N$ demonstrations with varying quality. Each demonstration $\mathcal{D}_{i}=\{\tau_{j}\}_{j=1}^{M_{i}}$ contains a set of $M_{i}$ trajectories sampled from a distinct, fixed policy $\pi_{i}$. Rather than considering $\mathcal{D}$ as a homogeneous set derived from one expert policy $\pi_{E}$, our approach focuses on leveraging the specific source of each demonstration to enhance the IRL framework. By identifying the varying qualities within this mixed collection of demonstrations, we aim to more accurately deduce the ground truth policy $\pi_{E}$, as opposed to a simplistic method that averages behaviors. To achieve this, it is essential to develop a model for the demonstrators $\pi_{i}$ that accurately reflects their suboptimal behavior relative to the optimal policy $\pi_{E}$.
A popular framework for modeling decision making is the Boltzmann rationality model, which is widely used in psychology (Baker, Saxe, and Tenenbaum 2009), economics (Luce 1959) and RL (Laidlaw and Dragan 2021). In this model, the likelihood of an agent choosing an action is proportional to the exponential of the reward associated with that action: $\pi(a|s;r,\beta)\propto\exp(\beta r(s,a))$, where the inverse temperature parameter $\beta$ controls the randomness of the choice. While this formulation provides an intuitive way to capture one form of suboptimality through $\beta$, it traditionally assumes that agents have unbiased knowledge about the true reward $r(s,a)$. To broaden this model for suboptimal demonstrations, we propose considering agents that follow a biased reward $r(s,a)+\epsilon(s,a)$, introducing $\epsilon$ as the deviation from the true reward. This modification enables us to represent an agent’s accuracy through the deviation $\epsilon$, adding a bias to their perceived rewards, and their precision through parameter $\beta$, adding variance to their action choices.
Within the context of the IRL framework described in Section 2, we recall that the expert policy $\pi_{E}$ aligns with the true reward, formulated as $r(s,a)=\theta^{\star\top}f(s,a)$. Applying the aforementioned concept, we can model demonstrator $i$’s perceived reward as a deviation from the true reward: $r_{i}(s,a)=(\theta^{\star}+\epsilon_{i})^{\top}f(s,a)$, where $\epsilon_{i}\in\mathbb{R}^{k}$. By incorporating this altered reward $r_{i}$ into the soft Bellman policy defined in Eq. (3), we derive a recursive parameterization for demonstrator $i$’s policy:
$\pi_{\theta^{\star},\epsilon_{i},\beta_{i}}(a_{t}|s_{t})\propto\exp(\beta_{i}(Q^{\textit{soft}}_{\theta^{\star}+\epsilon_{i}}(s_{t},a_{t}))),$ (6)
where $\beta_{i}\in[0,\infty)$, and the normalizing factor not shown above should be modified to account for $\beta_{i}$: $V^{\textit{soft}}_{\theta^{\star}+\epsilon_{i}}(s_{t})=\frac{1}{\beta_{i}}\log\sum_{a_{t}\in\mathcal{A}}\exp(\beta_{i}Q^{\textit{soft}}_{\theta^{\star}+\epsilon_{i}}(s_{t},a_{t}))$.
@@
It is important to note that this model extends outside the scope of problems where the true reward $r$ is an affine transformation of some feature vector $f$. In the general setting, we can parameterize demonstrator $i$’s reward as $r_{i}=r+\epsilon_{i}$, a combination of the true reward $r\in\mathcal{R}$ and their deviation $\epsilon_{i}\in\mathcal{R}$, where $r$ and $\epsilon_{i}$ can be represented by neural networks to allow for general function approximation. We opt to formulate our proposed model within the classic feature matching scope, as this facilitates direct comparison with the seminal works used to derive our results in the following section. However, our proposed model’s applicability extends beyond this traditional framework, as we elaborate in Section 4, and demonstrate through experiments in Section 5. Until then, we utilize the feature matching framework described, referencing it as standard IRL.